A company had total revenues of $114 million, operating profit margin of 22%, and depreciation and amortization expense of $11 million over the trailing twelve months. The company currently has $44 million in total debt and $11 million in cash and cash equivalents. If the company's market capitalization (market value of its equity) is $571 million, what is its EV/EBITDA ratio? Round to one decimal place.
Answer #1

EBIDTA = revenues*oper margin = 114*0.22 = 25.08

Enterprise value = Equity value+ MV of debt
- Cash & Cash Equivalents
Enterprise value = 571+44-11
Enterprise value = 604

EV/EBIDTA = 604/24.08 =

25.1